With the wind, the plane can fly 3000 miles in 5 hours. Against the same wind, the trip takes 6 hours. Find the airspeed of the plane.
Answer by stanbon(75887)

With the wind, the plane can fly 3000 miles in 5 hours. Against the same wind, the trip takes 6 hours. Find the airspeed of the plane.
---------------
with-wind DATA:
dist = 3000 miles ; time = 5 hrs ; rate = d/t = 3000/5 = 600 mph
-------------
against-wind DATA:
dist = 3000 miles ; time = 6 hrs ; rate = 3000/6 = 500 mph
----------------
Equations:
p + w = 600
p - w = 500
------------------
Add and solve for "p":
2p = 1100
p = 550 (speed of the plane in still air)
---
Solve for "w":
p + w = 600
550 + w = 600
w = 50 mph (speed of the wind)
==================================
Cheers,
Stan H.